[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu May 18, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu May 18 11:48:56 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
There were new NWPS issues last night that appear to be related to the run  
length in the GUI. Chris Mello and Craig will investigate.

Forecasters attempted to run the rainfall program for a disturbance near  
Haiti, but did not get any output back. However, Dave confirmed that the  
rainfall products are on the web. TAFB plans to run the rainfall program  
again this morning. We will check to see if there are any issues then.

There were no shift log entries on problems with the cumulus N-AWIPS  
workstation. Evelyn will be working on the system today.

HSU:

Other items:
GOES-16 data continues to be missing from N-AWIPS. Monica will look into  
this.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170518/6e4b0c64/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list